Table 8.1. Mass data for analysis of alum. Mass of
Answer #1
  1. % age SO42-by mass in dry alum

Molar mass of KAl(SO4)2 is 258.2050 g/mol

Sulfur

32.0

2 (no. of atoms)

Oxygen

15.9

8(no. of atoms)

Molar mass of (SO4)2 is 192.1252 g/mol

% age SO42-by mass in dry alum = Molar mass of (SO4)2-/ Molar mass of KAl(SO4)2 * 100

(192.1252/258.2050)*100 = 74.40%

% age SO42-by mass in wet alum

Molar mass of KAl(SO4)2*12H2O is 474.3884 g/mol

Molar mass of (SO4)2 is 192.1252 g/mol

% age SO42-by mass in wet alum = (Molar mass of (SO4)2-/ Molar mass of KAl(SO4)2 * 12 H2O )*100

192.1252/474.3884 = 40.499 %

So, we see that there is decrease in (SO4)2-% by mass in wet alum

2. The qualitative test for sulfate uses the insolubility of barium sulfate (BaSO4). When an aqueous solution of a barium salt (usually BaCl2) is mixed with an aqueous solution containing sulfate, a white precipitate of insoluble BaSO4 forms according to the net ionic equation: Ba2+ (aq) + SO42- (aq) → BaSO4 (s). A positive test for SO42-, therefore, is the observation of a white precipitate when an aqueous BaCl2 solution is mixed with the aqueous test solution.

As precipitate contains sulfate from BaSO4, Ba2+ and SO42- from Ba CO3 only Ba2+

Total SO42- = 192.1252 g/mol

Total weight of precipitate = Molar mass of BaSO4+ molar mass of Ba from BaCO3 = 192.12+137.32 = 329.44

% age SO42-by mass in precipitate = 192.12/329.44 = 58.3%

So, we see that there is decrease in (SO4)2-% by mass

3. The decanted liquid if was cloudy, that means some more crystals can be precipitated out of it. This second crop of crystals may also be collected, but if you choose to collect these crystals, they should be kept separate from the main crop. It is always good laboratory practice to keep different crops of crystals separate until the identity and purity of each crop is determined.

second crops almost always contain more impurities than the first crop and the time needed to purify them sometimes far outweighs the additional yield. There might be increase in the percent SO42- by mass